# P-NAL: an Effective and Interpretable Entity Alignment Method

### v1

@@ Abstract.
Entity alignment (EA) aims to find equivalent entities between two Knowledge Graphs. Existing embedding-based EA methods usually encode entities as embeddings, triples as embeddings’ constraint and learn to align the embeddings. The structural and side information are usually utilized via embedding propagation, aggregation or interaction. However, the details of the underlying logical inference steps among the alignment process are usually omitted, resulting in inadequate inference process. In this paper, we introduce P-NAL, an entity alignment method that captures two types of logical inference paths with Non-Axiomatic Logic (NAL). Type I is the bridge-like inference path between to-be-aligned entity pairs, consisting of two relation/attribute triples and a similarity sentence between the other two entities. Type II links the entity pair by their embeddings. P-NAL iteratively aligns entities and relations by integrating the conclusions of the inference paths. Moreover, our method is logically interpretable and extensible due to the expressiveness of NAL. Our proposed method is suitable for various EA settings. Experimental results show that our method outperforms state-of-the-art methods in terms of Hits@1, achieving 0.98+ on all three datasets of $DBP15K$ with both supervised and unsupervised settings. To our knowledge, we present the first in-depth analysis of entity alignment’s basic principles from a unified logical perspective.
1. Introduction
@@
Knowledge graphs (KGs), which store massive facts about the real world, is a fruitful attempt aiming to enhance semantic driven information processing ability. KGs can be used by various application domains, such as question answering, recommender systems and language representation learning (knowledge graph enhanced language model)  (Ji et al., 2021; Logan IV et al., 2019). The information contained in each individual KG project, such as DBpedia  (Auer et al., 2007) and YAGO  (Suchanek et al., 2007) is limited. So the task of entity alignment (EA) is proposed to increase KG completeness. The EA task consists of integrating two or more KGs into a same KG by aligning nodes that refer to the same entity.
@@
There are many embedding-based EA methods  (Fanourakis et al., 2023) that leverage deep learning techniques to represent entities with low-dimensional embeddings, and align entities with a similarity function on the embedding space. KGs’ triples and seed alignments are usually seen as embeddings’ constraint during the training process of such embedding model. The structural and side information of KGs are usually utilized via embedding propagation, aggregation or interaction. Generally speaking, there are some crucial shortcomings of embedding-based EA methods: 1. They lack complex reasoning capability. Some of them are enhanced by paths  (Cai et al., 2022), however, due to the nature of vector representation, it is not easy to perform or approximate symbolic reasoning on such paths. 2. They lack interpretability in the models, so they have to rely solely on numerical evaluation metrics to evaluate their performance. Thus the cons and pros of their model design may not be properly evaluated. 3. The absence of a unified framework explaining the mechanism of embedding learning and processing renders their semantic or structural learning capability quite mysterious.
@@
Apart from embedding-based methods, there exist a group of methods that directly estimates entity similarities from the contextual data (path) that are available in the two input KGs. We refer to them as “path-based” methods. Also, we refer to the estimation of entity similarities by processing and aggregating the paths as “similarity inference”. There is a potential advantage that path-based methods can capture fine-grained matches of neighbors while the traditional embedding-based methods can’t. There are also emerging methods that combines the idea of embedding learning and path reasoning. We coin the term “embedding-path” to refer to such methods. More recently, path-based (such as PARIS+  (Leone et al., 2022)) and embedding-path methods (such as BERT-INT  (Tang et al., 2020) and FGWEA  (Tang et al., 2023)) are starting to surpass the performance of traditional embedding-based methods. However, they failed to handle the similarity inference appropriately to some extent, possibly due to the lack of proper formalization of the inference paths and steps.
To address the aforementioned issues of embedding-based and embedding-path methods, we carefully examine the similarity inference of EA from the logical perspective. Thus we propose a path-based EA method P-NAL, where P stands for PARIS  (Suchanek et al., 2011) and NAL for Non-Axiomatic Logic  (Wang, 2013). PARIS is an unsupervised non-neural EA method with competitive performance on benchmark datasets  (Leone et al., 2022). NAL is a term logic with a specific semantic theory and its design suits KG tasks (see Section 2.3). P-NAL reinterprets and extends the traditional EA system PARIS with the help of NAL. We formalize the similarity inference as using NAL’s revision inference rule to aggregate the evidences of two types of inference paths. Type I is the bridge-like inference path between to-be-aligned entity pairs, consisting of two relation/attribute triples and a similarity between the other two entities. It can be seen as the fundamental alignment evidence (signal), and we coin the term “align-bridge” to refer to such type. Similarity inference of align-bridge consists of clarifying premises, performing several inference steps and obtaining conclusions. Type II is the direct path linking the to-be-aligned entity pairs by the representations of their name or description. We obtain such representations (embeddings) through the deep language model BERT. There are other two types of inference paths (which align relations), III and IV, that is not performed by our method but depicted for theoretical purpose.
P-NAL’s adopts an iterative aligning strategy, and for each iteration it first performs similarity inference, then it uses a matching technique (rBMat algorithm with modification, see Section 3.4) to obtain EA results. It also infers the matching of relations in each iteration. Although P-NAL is path-based, it does use embedding technique minimally, just to embed the literals in KGs (entity names/descriptions and attribute values). The name/description embeddings are utilized by path type II and the attribute value embeddings are utilized by path type I.
P-NAL is simple, highly interpretable and self-explanatory. The design of P-NAL has simple intuition. Although the overall implementation may seem a little complicated, P-NAL avoids using unnecessarily complicated mathematical objects and each step in it can be easily understood. It is interpretable in the sense that similarity inference and relation inference of P-NAL shares the same logical foundation and uses interpretable logical inference steps. P-NAL is self-explanatory in the sense that it generates a log file of evidences for the alignments after each iteration so we can inspect the log files. This feature enhances the troubleshooting capacity of us to some extent during the development process of P-NAL. For example, inspecting the faulty alignments in the evidence file inspired many decision choices in this paper. Also, P-NAL is extensible because NAL can express and process many different reasoning patterns and logical structures, so P-NAL can be extended to tackle other challenges in the EA process in future research.
@@
Experiments on cross-lingual EA dataset $DBP15k$ demonstrate that P-NAL outperforms 12 existing EA methods, including both supervised and unsupervised state-of-the-art approaches, in 5 different configuration groups. Ablation study shows that our design choices jointly boost the overall performance of our model. With competitive EA performance, we conjecture that P-NAL’s similarity inference captures the essence of current EA task (with attribute triples), which means that aligning by such paths is both intuitive and effective.
@@
To our knowledge, we present the first in-depth analysis of entity alignment’s basic principles from a unified logical perspective. Moreover, P-NAL is the first to integrate NAL in the EA task. P-NAL might also help explain the mechanism of other EA methods, as discussed in Section 4.
Outline. Section 2 gives a formal definition of EA and overviews the related works and relevant background knowledge of NAL. Section 3 first sketches our overall EA framework, then elaborates our design, including inference paths, literal utilization, matching algorithm, etc. Section 3.5 elaborates the overall structure of P-NAL. Section 4 discusses the relation with other methods. Our experimental results are presented in Section 5. We conclude in Section 6.
2. Preliminaries
2.1. Knowledge Graph and Entity Alignment
@@
Knowledge graphs (KGs) are knowledge bases that store knowledge in the form of so-called facts or triples. We refer to (head, relation, tail) and (head, attribute, literal) as relation and attribute triples, respectively. Examples of both triple types are (New_Zealand, capital, Wellington) and (New_Zealand, establishedDate, “1947-11-25”), respectively. The arguments head and tail represent entities, relation is a relationship that holds between two entities, and attribute is a special type of relation that holds between an entity and a literal. Entities can be seen as graph nodes and they usually denote real-world objects, while literals are used to identify values for strings, numbers or dates. To summarize, a KG is characterized with a number of relation triples from $\mathcal{E}\times\mathcal{R}\times\mathcal{E}$ and a number of attribute triples from ×$\mathcal{E}\times\mathcal{A}\times\mathcal{L}$ , where $\mathcal{E},\mathcal{R},\mathcal{A}$, and $\mathcal{L}$ indicate the set of entities, relations, attributes and literals, respectively.
@@
The entity alignment (EA) problem is typically defined between two KGs, $\mathcal{KG}_{1}$ and $\mathcal{KG}_{2}$, where the task consists of finding equivalences (so-called alignment) between the set of entities $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ of the two KGs. Sometimes there exists a set of given equivalences that can be used as supervision. This set $\mathcal{S}$ is known as seed alignment set. The supervised EA methods are allowed to utilize the information of $\mathcal{S}$ to infer the equivalences of other entities. We assume that there exists a ground truth set $\mathcal{G}=\{(e,e^{\prime})\in\mathcal{E}_{1}\times\mathcal{E}_{2}|\ e\equiv e^{\prime}\}$ that includes all known equivalences between pairs of entities. The ground truth set $\mathcal{G}$ is usually used to evaluate the performance of an EA method, by comparing it with the output alignment set of the method.
2.2. Related Work
@@
Generally speaking, there are three families of EA methods: embedding-based, path-based and embedding-path methods, as elaborated in this section.
In recent years, embedding-based methods have become mainstream for addressing the EA task  (Tang et al., 2023; Fanourakis et al., 2023). Their main idea is to embed the nodes (entities) and edges (relations or attributes) of a KG into a low-dimensional vector space that preserves their similarities in the original KG. Embedding-based EA methods usually consists of three parts: the embedding module, the alignment module and the matching module. For the embedding module, translational methods and graph neural network (GNN) methods are the most popular. Translational methods, such as MTransE  (Chen et al., 2016), usually optimize a margin-based loss function to learn the structural information (relation triples) of a KG. On the other hand, GNN methods recursively aggregate the representations of neighboring nodes with graph convolutional networks (GCNs) or graph attention networks (GATs). The representative ones are RDGCN  (Wu et al., 2019) and RREA  (Mao et al., 2020b), respectively. The alignment module maps the entity embeddings in different KGs into a unified space. There are generally three techniques  (Fanourakis et al., 2023) for this module: 1. Sharing the embedding space by using the margin-based loss to enforce the seed alignment entities’ embeddings from different KGs to be close. 2. Swapping the triples of seed alignment entities. 3. Mapping the entity vectors from one embedding space to the other using a transformation matrix. The matching module generates the final alignment result. Common practices use the cosine similarity, the Manhattan distance, or the Euclidean distance between entity embeddings to measure their similarities and then performs a specific matching algorithm based on the similarity scores.
@@
Apart from embedding-based methods, there exist a group of methods that directly estimates entity similarities from the contextual data (path) that are available in the two input KGs. We refer to them as “path-based” methods. There is a potential advantage that path-based methods can capture fine-grained matches of neighbors while the traditional embedding-based methods can’t. Embedding-based methods may suffer from the negative influence from the dissimilar neighbors, according to  (Tang et al., 2020). The distinction between embedding-based and path-based methods is sometimes obscure.
@@
There are also emerging “embedding-path” methods that combines the idea of embedding learning and path reasoning. More recently, path-based and embedding-path methods are starting to surpass the performance of traditional embedding-based methods. Our proposed method P-NAL is a path-based method that uses embedding minimally, so in the following part we introduce several advanced path-based and embedding-path methods. Moreover, these methods are selected for comparison with P-NAL in our experiments.
PARIS  (Suchanek et al., 2011) is a classic unsupervised non-neural EA method with competitive performance on benchmark datasets  (Leone et al., 2022). It is purely path-based. PARIS introduces the concept “functionality” into the field of EA to enhance the validity of similarity inference paths. Functionality generally corresponds to the uniqueness of related things, for example a man can only have one father but multiple friends, so $fun(father)$ is close to 1 and $fun(friend)$ is relatively lower, where $fun()$ represents functionality of a relation or attribute. See  (Suchanek et al., 2011) for more details about functionality. With functionality, PARIS constructs a probabilistic model that estimates the probabilities of an entity $x$ in $\mathcal{KG}_{1}$ being equivalent to another entity $x^{\prime}$ in $\mathcal{KG}_{2}$. Here is the formula for $Pr\left(x\equiv x^{\prime}\right)$:
$1-\prod_{r(x,y),r^{\prime}(x^{\prime},y^{\prime})}\left(1-Pr\left(r\subset r^{\prime}\right)\times fun(r^{-1})\times Pr\left(y\equiv y^{\prime}\right)\right)$
As depicted in the above formula, PARIS estimates the equivalence probabilities by integrating paths that connects corresponding entities. It also find subrelations between the two ontologies of KG with the following equation. Subrelations, such as $r\subset r^{\prime}$, intuitively means a correspondence of two relations of different KGs such that one relational fact of $r$ in $\mathcal{KG}_{1}$ implies the existence of a corresponding relational fact of $r^{\prime}$ in $\mathcal{KG}_{2}$. Here is the formula for $Pr\left(r\subset r^{\prime}\right)$:
$\frac{\Sigma_{r(x,y)}\left(1-\prod_{r^{\prime}(x^{\prime},y^{\prime})}\left(1-Pr\left(x\equiv x^{\prime}\right)\times Pr\left(y\equiv y^{\prime}\right)\right)\right)}{\Sigma_{r(x,y)}\left(1-\prod_{x^{\prime},y^{\prime}}\left(1-Pr\left(x\equiv x^{\prime}\right)\times Pr\left(y\equiv y^{\prime}\right)\right)\right)}$
With the help of subrelations’ measurement, PARIS generalizes the equation of $Pr\left(x\equiv x^{\prime}\right)$ to the case where the two ontologies do not share common relations. Therefore, PARIS recursively aligns the entities and the equivalence probability of $x\equiv x^{\prime}$ depends recursively on other equivalence probabilities. In each iteration, the probabilities are re-calculated based on the equivalences and subrelations of the previous iteration. Initial equivalences are computed between attribute literals based on a certain string distance measurement.
PARIS+  (Leone et al., 2022) is a variant of PARIS that makes a simple refinement and works in the absence of attribute triples. It processes the seed alignment information to generate synthetic attribute triples. That is, for every pair of seed alignments ($x$, $x^{\prime}$), it creates the attribute triples ($x$, EA:label, string($x$)) and ($x^{\prime}$, EA:label, string($x$)), where EA:label is a synthetic relation. Thus, the reverse of the relation EA:label is designed to be highly functional in order to let the model match the seed alignments easily. P-NAL adopts the same refinement as PARIS+.
BERT-INT  (Tang et al., 2020), an embedding-path EA method, uses the well-known transformer model BERT to embed the entities and literals. It calculates the cosine similarity of the entity name/description embedding. Then it proposes an interaction model that compares each pair of neighbors or attributes (which forms a path from the source entity to the target entity) to obtain the neighbor/attribute similarity score. The name/description similarity vector, neighbor similarity vector and attribute similarity vector are concatenated and applied to a MLP layer to get the final similarity score.
FGWEA  (Tang et al., 2023) is a three-step progressive optimization algorithm for EA and it can be classified as an embedding-path EA method. First, the entity names and concatenated attribute triples are used for semantic embedding matching to obtain initial anchors. Then in order to approximate GWD (Gromov-Wasserstein Distance  (Peyré et al., 2016)), FGWEA computes cross-KG structural and relational similarities, which are then used for iterative multi-view optimal transport alignment. Finally, the Bregman Proximal Gradient algorithm  (Xu et al., 2019) is employed to refine the GWD’s coupling matrix.
There are also a few works that focus on the interpretability or explanability of EA, such as LightEA  (Mao et al., 2022) and ExEA  (Tian et al., 2023). LightEA is an interpretable non-neural EA method. It is inspired by a classical graph algorithm, label propagation  (Zhu and Zoubin, 2002). First, it generates a random orthogonal label for each seed alignment entity pair. Then, the labels of entities and relations are propagated according to the three views of adjacency tensor. Finally, LightEA utilizes sparse sinkhorn iteration to address the assignment problem of alignment results.
The ExEA framework, proposed by  (Tian et al., 2023), aims to explain the results of embedding-based EA. It generates semantic matching subgraphs as explanation by matching semantically consistent triples around the two aligned entities. ExEA devises an alignment dependency graph structure to gain deeper insights into the explanation.
The recent literature of EA is abundant, focusing on many different aspects or procedures of entity alignment apart from the aforementioned ones, such as utilizing attribute triples  (Liu et al., 2020; Sun et al., 2017), utilizing literals  (Gesese et al., 2021; Chen et al., 2018) , sample mining  (Liu et al., 2022; Mao et al., 2021a), reinforcement learning  (Guo et al., 2022), matching algorithm  (Lin et al., 2023; Dao et al., 2023; Mao et al., 2021b; Xu et al., 2020; Zeng et al., 2020), iterative strategy  (Liu et al., 2023; Mao et al., 2020a) and unsupervised learning  (Jiang et al., 2023b, a; Liu et al., 2022; Luo and Yu, 2022; Zhao et al., 2022). There are also some surveys for EA  (Fanourakis et al., 2023; Zeng et al., 2021; Sun et al., 2020; Mao et al., 2022). Besides graph structural, attribute and literal information, there are other information forms researched by the EA community, such as temporal, spatial and graphical information, however, these topics are beyond the scope of this paper.
2.3. A Brief Introduction to NAL
NAL (Non-Axiomatic Logic)  (Wang, 2013) is a logic designed for the creation of general-purpose AI systems, by formulating the fundamental regularities of human thinking in a general level. It can be used as the logical foundation of a (non-axiomatic) inference system. Traditional inference systems are usually based on model-theoretic semantics, while under the assumption of insufficient knowledge and resources, NAL is a term logic basing on experience-grounded semantics  (Wang, 2005). The meaning of a term in NAL, to the inference system, is determined by its role in the experience (which will be explained later), that is, how it has been related to other terms in the past. The truth-value of a statement in NAL is determined by how it has been supported or refuted by other statements in the past.
In this paper we only utilize a fraction of NAL’s syntax and inference capability (for EA). We will now introduce the relevant parts of its syntax. A term in NAL can either be atomic or compound. An atomic term is a word (string) or a variable term. Independent variable, such as “#$\#x$”, represents any unspecified term under a given restriction, and intuitively correspond to the universally quantified variable in first-order predicate logic. Dependent variable, such as “\$$\$y$”, represents a certain unspecified term under a given restriction, and intuitively correspond to the existentially quantified variable. A compound term consists of term connector and components (which are themselves terms). A basic statement has the form of “subject copula predicate”, where subject and predicate are terms. There are multiple types of copula and each type has a corresponding statement type, including: 1.Inheritance (“$A\rightarrow B$”, where $A$ and $B$ are terms) which intuitively means “B is a general case of A”; 2.Similarity (“$A\leftrightarrow B$”) which intuitively means “A is similar with B”; 3.Implication, which is a higher-order copula (“$P\Rightarrow Q$”, where $P$ and $Q$ are statements), intuitively means “P implies Q” (different from the “material implication”, it requires $P$ to be related to $Q$ in content because NAL is a term logic that uses syllogistic inference rules and only derives conclusions that are related in content). A sentence is a statement together with its truth-value. An intensional set with only one component, for example, “$[red]$” intuitively means “red things”. Term connector*” (product) combines multiple component terms into an ordered compound term such as $\left(*,A,B\right)$, which intuitively means “an anonymous relation between A and B”. Compound terms are usually written in the prefix format, that is the term connector is written in the first place. Statement $\wedge$” can be seen as the conjunction operator of propositional logic.
NAL is “non-axiomatic” in the sense that the truth-value of a conclusion in the inference system does not indicate how much the conclusion agrees with the “state of affair” in the world, or with a constant set of assumptions (the axioms), but how much it is supported by the evidence provided by the past experience of the system. Experience means the inference system’s history of interaction with the environment or equivalently the input sentences. The acquisition of experience may involve sensorimotor mechanism and sensation-perception process, which is beyond our scope. The information source of a sentence is characterized as its evidence. The inference rules of NAL coherently pass on the evidential information from the premises to the conclusion, so the premises can be seen as the evidence of the conclusion. The input sentences can be seen as a synthesis of virtual positive and negative evidences. Assume the available amount of positive evidence and negative evidence of a statement are written as $w^{+}$ and $w^{-}$, respectively, then the total amount of evidence is $w=w^{+}+w^{-}$. The frequency of the statement is $f=w^{+}/w$, and the confidence of the statement is $c=w/(w+k)$, where $k$ is a positive constant representing “evidential horizon”. We take $k$ = 1 in our implementation. Frequency intuitively means “the degree of truth” and confidence intuitively represents “the total amount of evidences”. The more evidences that the statement have considered, the higher confidence value. The truth-value attached to the statement is the ordered pair $<f,c>$ and it is often written right after the statement.
NAL uses syllogistic (rather than truth-functional) inference rules. Among them the revision rule merges evidences for the same statement collected from different sources together, so it can settle inconsistency among the system’s sentences. It is very useful in our approach. The relevant rules with corresponding truth functions are all listed in Table 1. Note that the inference rules are not domain-specific. There are three extended boolean operators  (Wang, 2013) in the calculation of truth functions:

### v2

# NALA: an Effective and Interpretable Entity Alignment Method

@@ Abstract
Entity alignment (EA) aims to find equivalent entities between two Knowledge Graphs. Existing embedding-based EA methods usually encode entities as embeddings, triples as embeddings’ constraint and learn to align the embeddings. However, the details of the underlying logical inference steps among the alignment process are usually omitted, resulting in inadequate inference process. In this paper, we introduce NALA, an entity alignment method that captures three types of logical inference paths with Non-Axiomatic Logic (NAL). Type I&II align the entity pairs and type III aligns relations. NALA iteratively aligns entities and relations by integrating the conclusions of the inference paths. Our method is logically interpretable and extensible by introducing NAL, and thus suitable for various EA settings. Experimental results show that NALA outperforms state-of-the-art methods in terms of Hits@1, achieving 0.98+ on all three datasets of DBP15K with both supervised and unsupervised settings. We offer a pioneering in-depth analysis of the fundamental principles of entity alignment, approaching the subject from a unified and logical perspective. Our code is available at https://github.com/13998151318/NALA.
{CJK*}
UTF8gbsn
NALA: an Effective and Interpretable Entity Alignment Method
Chuanhao Xu, Jingwei Chengthanks: Corresponding author, Fu Zhang School of Computer Science and Engineering, Northeastern University, China 2201892@stu.neu.edu.cn, {chengjingwei, zhangfu}@mail.neu.edu.cn
1 Introduction
Figure 1: An overview illustration of NALA.
@@ 1 Introduction
Knowledge graphs (KGs), which store massive facts about the real world, expresses massive information in a form closer to human cognition. KGs can be used by various application domains, such as question answering, recommender systems and language representation learning (knowledge graph enhanced language model)  Ji et al. (2021); Logan IV et al. (2019). The information contained in each individual KG project, such as DBpedia  Auer et al. (2007) and YAGO  Suchanek et al. (2007) is limited. So the task of entity alignment (EA) is proposed to increase KG completeness. The EA task consists of integrating two or more KGs into a same KG by aligning nodes that refer to the same entity.
@@
There are many embedding-based EA methods  Fanourakis et al. (2023) that leverage deep learning techniques to represent entities with low-dimensional embeddings, and align entities with a similarity function on the embedding space. KGs’ triples and seed alignments are usually seen as embeddings’ constraint during the training process of such embedding model. The structural and side information of KGs are usually utilized via embedding propagation, aggregation or interaction. Generally speaking, there are some crucial shortcomings of embedding-based EA methods: First, they possibly lack complex reasoning capability. Some of them are enhanced by paths  Cai et al. (2022), however, due to the nature of vector representation, it is not easy to perform or approximate symbolic reasoning on such paths. Second, they lack interpretability in the models, so they have to rely solely on numerical evaluation metrics to evaluate their performance. Thus the cons and pros of their model design may not be properly evaluated. Third, the absence of a unified framework explaining the mechanism of embedding learning and processing renders their semantic or structural learning capability quite mysterious.
@@
Apart from embedding-based methods, path-based methods directly estimates entity similarities from the contextual data (path) that are available in the two input KGs. A "path" usually refers to an interconnected sequence of edges that links two entities of different KGs. The edges can be either relations or entity similarities. We refer to the estimation of entity similarities by processing and aggregating the paths as "similarity inference". There is a potential advantage that path-based methods can capture fine-grained matches of neighbors while the traditional embedding-based methods can’t. There are also emerging methods that combine the idea of embedding learning and path reasoning. More recently, path-based (such as PARIS+  Leone et al. (2022)) and combined methods (such as BERT-INT  Tang et al. (2020) and FGWEA  Tang et al. (2023)) are starting to surpass the performance of traditional embedding-based methods. However, they failed to handle the similarity inference appropriately to some extent, possibly due to the lack of proper formalization of the inference paths and steps.
To address the aforementioned issues of existing methods, we carefully examine the similarity inference of EA from the logical perspective. Thus we propose a path-based EA method NALA, where NAL stands for Non-Axiomatic Logic  Wang (2013) and "A" for align. NAL is a term logic with a specific semantic theory and its design suits KG tasks (see Section 2.3).
As illustrated in Figure 1, NALA adopts an iterative entity alignment strategy with two modules, namely similarity inference module and matching module. For each iteration, it first performs the similarity inference module. The module has three functions: 1) Search for three types of paths across the KGs. 2) Inference on the path instances with fixed inference rules defined by NAL. 3) Aggregate the conclusions of the paths. Type I&II paths yield conclusions on similarities of entity pairs. Type III paths yield conclusions on substitutability (or, inheritance) of relation or attribute pairs. We use BERT embedding to assist path inference by obtaining similarity among entity names and attribute values, which constitutes some premises of the paths. Then NALA uses the matching module to obtain 1-to-1 EA results of the current iteration. We propose an algorithm, namely rBMat algorithm with swapping step (see Section 3.2) for the matching module.
@@
Experiments on cross-lingual EA dataset DBP15K demonstrate that NALA outperforms SOTA EA methods in 5 different setting groups (including both supervised and unsupervised scenarios), showcasing the effectiveness of our proposed logical similarity inference module and matching module. Ablation study shows that our design choices jointly boost the overall performance of NALA.
Our contributions can be summarized as:
  • We propose an interpretable EA framework NALA, which tackle the EA problem with similarity inference phase and matching phase. Various types of logical paths are formalized within the similarity inference phase.
  • NALA aligns entities and relations simultaneously with a unified yet extensible logical framework.
  • Our framework bridges the gap between embedding-based and path-based EA.
  • Our proposed method achieves SOTA on a widely used EA dataset DBP15K’s various settings.
@@
  • We present the first in-depth analysis of EA’s basic principles from a unified logical perspective, and help explain the mechanism of other EA methods.
2 Preliminaries
2.1 Knowledge Graph and Entity Alignment
@@
KGs. Knowledge graphs (KGs) are knowledge bases that store knowledge in the form of triples (or "facts"). We refer to (head, relation, tail) and (head, attribute, literal) as relation and attribute triples, respectively. Examples of both triple types are (New_Zealand, capital, Wellington) and (New_Zealand, establishedDate, "1947-11-25"), respectively. To summarize, a KG is characterized with a number of relation triples from $\mathcal{E}\times\mathcal{R}\times\mathcal{E}$ and a number of attribute triples from ×$\mathcal{E}\times\mathcal{A}\times\mathcal{L}$ , where $\mathcal{E},\mathcal{R},\mathcal{A}$, and $\mathcal{L}$ indicate the set of entities, relations, attributes and literals, respectively.
@@
EA. The entity alignment (EA) problem is typically defined between two KGs, $\mathcal{KG}_{1}$ and $\mathcal{KG}_{2}$, where the task consists of finding equivalences (so-called alignment) between the set of entities $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ of the two KGs. Sometimes there exists a set of given equivalences that can be used as supervision. This set $\mathcal{S}$ is known as seed alignment set. We assume that there exists a ground truth set $\mathcal{G}=\{(x_{1},x_{2})\in\mathcal{E}_{1}\times\mathcal{E}_{2}|\ x_{1}\equiv x_{2}\}$ that includes all known equivalences between pairs of entities. We use the ground truth set to evaluate the performance of our method. We use the subscript of an entity identifier ($x$ or $y$) or a relation identifier ($r$) to represent which KG it comes from.
2.2 Represent KGs with NAL
A brief introduction to NAL is presented in Appendix A.
In this paper, every entity, literal or relation is regarded as an atomic term in NAL. Triple ($x$$r$$y$) is reinterpreted as inheritance statement (*, $x$$y$) $\rightarrow$ $r$. Its intuitive meaning is "The relation between $x$ and $y$ is a specialization of relational term $r$". The triples (or "facts") of the KGs can be seen as absolutely true (for frequency) and with sufficient evidence (for confidence) to some extent, so the truth-value attached to the statement is $\left\langle 1,1\right\rangle$. Entity equivalency $\ x_{1}\equiv x_{2}$ can be seen as an extreme case of entity similarity $x_{1}$leftrightarrow  $x_{2}$, so we align entities by similarity inference. As for relations, the inheritance statement $r_{1}$rightarrow $r_{2}$ intuitively represents a correspondence of two relations of different KGs such that one relational fact of $r_{1}$ in $\mathcal{KG}_{1}$ implies the existence of a corresponding relational fact of $r_{2}$ in $\mathcal{KG}_{2}$.
We automatically duplicates every original KG triple ($x$$r$$y$) with a reversed triple ($y$$r^{-1}$$x$) upon KG loading, where $r^{-1}$ represents the reverse relation or attribute of $r$.
Inference path. We define an instance of inference path as a premise set of NAL sentences (triples, similarities, etc.) and a series of corresponding inference steps which will eventually lead to a conclusion sentence. The premise sentences are either in the KGs or inferred from the KGs. A type of inference path is a shared form of paths and it can be instantiated with concrete entities and relations. It is usually utilized for a certain purpose, such as aligning entities or aligning relations.
2.3 Why NAL
Actually there might be many different logical systems that are qualified to represent the similarity inference process of EA. However, we believe that the non-axiomatic nature of NAL fits in the domain of knowledge graph better than those axiomatic logical systems, because real world KGs need to deal with the problem of open-domain and alterable, incomplete or conflicting facts. Fundamentally, the tasks of knowledge graph (such as EA), fits well with the assumption of insufficient knowledge and resources  Wang (2013), which is the basic assumption of NAL.
2.4 Related Work of EA
@@
Generally speaking, there are three families of EA methods: embedding-based, path-based and combined methods, as elaborated in this section.
In recent years, embedding-based methods have become mainstream for addressing the EA task  Tang et al. (2023); Fanourakis et al. (2023). Their main idea is to embed the nodes (entities) and edges (relations or attributes) of a KG into a low-dimensional vector space that preserves their similarities in the original KG. Embedding-based methods may suffer from the negative influence from the dissimilar neighbors, according to  Tang et al. (2020).
@@
In addition to embedding-based methods, there exist path-based methods that directly estimates entity similarities from the contextual data (path) that are available in the two input KGs. The distinction between embedding-based and path-based methods is sometimes obscure.
@@
There are also emerging methods that combine the idea of embedding learning and path reasoning. More recently, path-based and combined methods are starting to surpass the performance of traditional embedding-based methods.
Our proposed method NALA inherits and develops the ideas of two path-based methods PARIS  Suchanek et al. (2011) and PARIS+. The two methods as well as some other EA methods that will be compared with our results are introduced in Appendix B.
3 The Proposed Method
The overall structure of NALA (as illustrated in Figure 1 and Algorithm 3) adopts an iterative aligning strategy, and for each iteration it first performs similarity inference, then it uses the matching module (rBMat algorithm with swapping step in Section 3.2) to obtain EA results. The inference within each iteration benefits from the alignment results (both entities and relations) of the previous iteration.